[figure: 9.7 cm]

MAGNETIC AND MÖSSBAUER STUDIES OF Bi_{m+1}Ti_{m−x−y}Fe_{x}X_{y}O_{3m+3} (X= Sn, Nb)

G.D. Sultanov, Z.Yu. Seidov, N.G. Guseinov Institute of Physics Academy of Sciences of Azerbaijan, Huseyn Javid St.33, 370143 Baku, Azerbaijan

We present the results of Mössbauer and magnetic studies of the polycrystalline compounds Bi_{m+1}Ti_{m−x−y}Fe_{x}X_{y}O_{3m+3} (X= Sn, Nb) with perovskite-like layered- type structure , where m is a number of layers in a perovskite-like package. The negative values of θ indicate that these compounds have an antiferromagnetic type interactions. In wide temperature range the Mössbauer spectra show features characteristic for relaxation effects. This indicates the of existence spin fluctuations.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : Z.Yu.Seidov

Address for correspondence : Institute of Physics Academy of Sciences of Azerbaijan, Huseyn Javid St.33, 370143 Baku, Azerbaijan

Email address : seidov@lan.ab.az